solve for x: (8+x)^-3=64. Do I raise both sides to the -3rd power?
Answer by nabla(475)

First of all I will change the negative exponent as follows:
(1/(8+x))^3=64
Now take the cube root of each side.
1/(8+x)=4
Which is the same as:
1/4=8+x
-31/4=x